OpenStudy (anonymous):

Using synthetic division, what is the quotient: (2x3 - 2x - 12) ÷ (x - 2) i will reward medal

OpenStudy (anonymous):

it is almost impossible to write synthetic division here but i will try first list the coefficients 2 0 -2 -12 don't forget the \(0\) for the missing \(x^2\) term

OpenStudy (anonymous):

then since you are dividing by \(x-2\) put a \(2\) on the side 2 0 -2 -12 2 ____________________

OpenStudy (anonymous):

ok i won't

OpenStudy (anonymous):

bring down the top 2 2 0 -2 -12 2 ____________________ 2

OpenStudy (anonymous):

then 2 times 2 is 4 2 0 -2 -12 2 4 ____________________ 2

OpenStudy (anonymous):

0 + 4 = 4 2 0 -2 -12 2 4 ____________________ 2 4

OpenStudy (anonymous):

ok i did that and everytime i got a number on the bottom i multiplied it by 2 and i got a remainder of 0.am i right?

OpenStudy (anonymous):

then 2 times 4 = 8 2 0 -2 -12 2 4 8 ____________________ 2 4

OpenStudy (anonymous):

-2 + 8 = 6 2 0 -2 -12 2 4 8 ____________________ 2 4 6

OpenStudy (anonymous):

yes, the remainder is 0

OpenStudy (anonymous):

2 0 -2 -12 2 4 8 12 ____________________ 2 4 6 0

OpenStudy (anonymous):

your quotient is \(2x^2+4x+6\) and the remainder is \(0\)

OpenStudy (anonymous):

that tells you \[2x^3 - 2x - 12=(x-2)(2x^2+4x+6)\]

OpenStudy (anonymous):

so the answer is\[(x-2)(x ^{2}+4x+6)\]

OpenStudy (anonymous):

well the actual answer to the division is \(2x^2+4x+6\)

OpenStudy (phi):

the quotient is only 2x^2 +4x+6

OpenStudy (anonymous):

ok thanks